# Redefining Fitness: Evolution as a Dynamic Learning Process

### v1

@@ Results
The polynomial form of the log-fitness in $p$ is familiar and simple to analyze. A distinction emerges depending on the relative magnitude of $b$ and $c$, Fig. 2A. When $b>c$ (or $WP>L^{2}$) the situation is simple because the log-fitness is always a growing function of $p$, with a single maximum at $p=1$ corresponding to full cooperation. Whatever the initial state of the population, cooperation will always emerge as the dominant strategy and will eventually fixate. Relative to the statistically independent case, cooperation now emerges from the fact that the population is effectively playing itself, so that increases in mutual exploitation lead to overall lower fitness. The situation is more interesting when $c>b$, because this leads to a fitness minimum at $p_{*}=\frac{c-b}{d}<1$. The existence of this fitness minimum creates a local fitness maximum at $p=0$ (mutual defection), which is locally stable. In these circumstances, sufficient cooperation $p>p_{*}$ must be present for the population to converge to the global fitness maximum of full cooperation ($p=1$). This corresponds to a simple condition for cooperation to emerge, related to Hamilton’s rule [54]. From Eq. 11, we identify benefits of cooperation as $B=b+dp$ and costs with $C=c$. Then $B>C$ is a condition for full cooperation to emerge, which is equivalent to $p>p_{*}$. The population dynamics around $p=p_{*}$ is characterized by a first-order transition – a "tipping point"– because there is a valley of fitness to be crossed to reach the global maximum, starting with generalized defection. This can be done locally, over space or on networks, as many authors have noted based on numerical experiments [27, 55]. We will return to this point below.
@@
Third, the situation is both more interesting and more complex when the agents’ actions become statistically dependent. In this case, there are three parameters at play in the joint distribution: four joint probabilities minus the normalization. We can parameterize this space in analogy to task selection as $p_{2}=f_{1|0}(1-p_{1})+f_{1|1}p_{1}$ and $f(s_{1}=1,s_{2}=1)=f_{1|1}p_{1}$, with the shorthand $f_{1|1}=f(s_{2}=1|s_{1}=1)$, $f_{1|0}=f(s_{2}=1|s_{1}=0)$. Note that $r=f_{1|1}-f_{1|0}$ is the regression of $p_{2}$ on $p_{1}$, which vanishes when the two agents are statistically independent. Thus, $r$ is a measure of how (co-)related the two states are. The log-fitness becomes
@@
Because this is linear in $p_{1}$ the nature of the maximum depends only on the sign of the term in the square brackets. This is a second version of Hamilton’s rule closer to the way Queller [22] expressed it. If we identify the benefits of interaction $B=b$, and costs $C=c$ we now obtain $rB>C$ as the condition for cooperation to emerge. Adding the $d$ term facilitates the emergence of cooperation via what Queller called a synergism coefficient, as $rB+df_{1|1}>c$. This transition has a geometric interpretation in the reciprocity-exploitation ($f_{1|1},f_{1|0}$) plane. We can write the cooperation condition as a line, $y=f_{1|1}$, $x=f_{1|0}$, $y=y_{0}+\beta x$, with $y_{0}=\frac{c}{b+d}$, $\beta=\frac{b}{b+d}$ , as shown in Figure 2C. Cooperation, $p_{1}\rightarrow 1$, emerges for values above this line. To be above the line requires controlling exploitation, optimized when $f_{1|0}=f(s_{2}=1|s_{1}=0)\rightarrow 0$ (which implies that punishment or retaliation, $f(s_{2}=0|s_{1}=0)\rightarrow 1$. It also requires reciprocity meaning that $f_{1|1}=f(s_{2}=1|s_{1}=1)\rightarrow 1$ ( so that one becomes safe from exploitation, $f(s_{2}=0|s_{1}=1)\rightarrow 0$). Note that the first of these conditions may take the population towards mutual defection  [56], which is suboptimal. The second is therefore necessary to reach the highest fitness globally. Below the line, cooperation cannot emerge. These outcomes agree with established results on the effects of punishment and rewards on cooperation [57, 58]. That is, both punishment and reciprocity promote the establishment of cooperation but alone punishment does not suffice [59]. The information in the population is maximal as agents coordinate but the amount of information at stake converges to zero at full cooperation because then there is no uncertainty left to explain.
@@
Figure 2: Phase diagram for the Prisoner’s Dilemma and the environmentally mixed game. A: Log-fitness versus cooperation probability $p$ for the well-mixed population. Note the fitness valley for $b<c$. Conditional dependent case conditions for cooperation. B: Probability of reciprocity $f_{1|1}$ versus environment, $q$. C: Probability of reciprocity versus exploitation $f_{1|0}$, for various $q$. Cooperation emerges in each case for parameters above the lines. See also Fig.S1
@@
$\ln w_{s_{1}}=a-c(e)s_{1}+b(e)s_{2}+ds_{1}s_{2},\quad{\rm with}\quad c(e)=c(1-e)-be,\ b(e)=b(1-e)-ce.$ (13)
Costs and benefits swap between the PD and the ToC, for which $\ln w_{s_{1}}=a+bs_{1}-cs_{2}+ds_{1}s_{2}$. If the two agents act independently, then the agent should always cooperate in the ToC because both $b,d>0$. If the two agents act independently but share the same probability $p_{1}=p_{2}=p$, the log-fitness is independent of $e$ because there is a symmetry between the two agents so that the benefits and costs of cooperation are fully socialized. Thus, in the fully mixed population, changes in the environment have no effect on the average population dynamics and the phase diagram is the same as in Fig.2B.
When the two agents coordinate, the situation becomes more interesting. If the agents can tell which game they are playing at each time, their probability of cooperation will be the selected strategy for each game weighed by their relative probability $q$. The other situation we consider is when agents act conditionally on each other independently of the state of the environment. Then we obtain the same condition for cooperation to emerge as in PD, but with costs and benefits that depend on the environment as in Eq. 13. Fig. 2B-C shows that when $q=0$, we obtain the PD condition (black line) in Fig. 2. As $q$ increases, both the ordinate at the origin and the slope decrease. At $q=1/2$ the slope is zero. Eventually, at $q=2/3$, ToC is sufficiently dominant and cooperation becomes unconditional.
This analysis is general and can be obtained for any combination of games leading to complex and interesting phase spaces, Fig. 2. As a result, we obtain transitions between different (optimal) population states as functions of environment probabilities, which play the role of control parameters, analogous to temperature in physical systems. We have therefore shown how mixing the prisoner’s dilemma with a cooperative game increases the selection for cooperation in non-trivial ways. This is likely a more realistic situation, as organisms and especially people, have a diversity of modes of interaction some competitive and some cooperative. Thus, the presence of explicitly cooperative channels some of the time can help fix unstable cooperation all of the time.
@@
The probabilistic interpretation of fitness readily extends to structured populations. Within a population, group identities specify additional layers of organization that condition phenotypes and therefore may contribute to individual selection.
We write the fitness of individual $i$ in group $g$ as $w_{gi}(g^{\prime},j)=Ap(\{s_{g^{\prime}j}\}|s_{gi})$, where $\{s_{g^{\prime}j}\}$ denotes the joint state of all neighbors $j,g^{\prime}$. Typically, these neighbors are not statistically independent from each other leading to synergistic interactions[23, 60]. This is important for finding clusters of states (genes, phenotypes) that collectively encapsulate adaptive functions [61].
However, the most common situation arises when interactions are dyadic (between pairs, only), so that the game occurs on a simple graph [62, 63, 35, 64]. In the log-fitness, this approximation is obtained when all the $s_{jg^{\prime}}$ are conditionally independent given $s_{ig}$ such that
$w_{gi}=\Pi_{j,g^{\prime}}Ap(s_{g^{\prime}j}|s_{gi}),\rightarrow\ln w_{g,i}=\ln A+\sum_{g^{\prime},j}\ln p(s_{g^{\prime}j}|s_{gi}).$ (14)
With these assumptions, we can write the log-fitness to be identical to Eq. 10, where the second player is replaced with the average over all connections, $s_{2}=1/K\sum_{g^{\prime},j}{s_{g^{\prime}j}}$. All our previous results still apply but require attention to group dependencies (see Supplementary Information for details).
@@
In particular, the agent should still defect unconditionally when it is independent from the other agents. When their probabilities coincide, in the well-mixed population $p_{gi}=p_{2}=p$, Eq. 11 still applies and a phase transition occurs at $p=p_{*}$, if $b<c$. When the states are not independent, the log-fitness is written compactly in terms of contributions from each interaction, weighed by the size of the population $1/K$,
@@
This equation is similar to what we obtained above, but with population definitions of $r$, $f_{1|1g},f_{1|0g}$, where the additional $g^{\prime},g$ labels denote conditional dependence across groups. The absence of the label means average across groups. The interesting possibility is that the agents’ behaviors are now different within and across groups in the sense that reciprocity $f_{1g^{\prime}|1g}<f_{1g|1g}$, and exploitation $f_{1g^{\prime}|0g}>f_{1g|0g}$, for $g^{\prime}\neq g$. Treating out group interactions as statistically independent leads to
$\ln w_{gi}=a+\underbrace{k_{g}\left(br_{g}+df_{1g|1g}-c\right)}_{\rm in-group}p_{gi}+\underbrace{\sum_{g^{\prime}\neq g}k_{g^{\prime}}\big{[}\left(dp_{g^{\prime}}-c\right)p_{gi}+bp_{g^{\prime}}\big{]}}_{\rm out-groups}.$ (16)
@@
This collapse of cooperation under external pressure can be averted in two ways. First, the agent may cease to interact with out-group members so that $k_{g^{\prime}\neq g}=0$, but this foregoes fitness benefits from the possible altruism of a large population. Fig.3A shows when these benefits promote cooperation, solving $rB>C$ for $k_{g}/K$ as a function of $p_{\bar{g}}$. As before, the contours define a boundary, above which more cooperation increases fitness. We also plot this solution in Fig3B for $f_{1g|1g}$ and $p_{\bar{g}}$, which can be compared to Fig.2B when $q=0$.
Second, cooperation or defection could conditionally depend on the group identity of the interaction partner. Then the in-group and out-group probabilities of cooperation $p_{gi|g^{\prime}}$ adapt differentially so that $p(gi|g)\rightarrow 1$, and $p(gi|g^{\prime}\neq g)\rightarrow 0$. This may still mean that the average cooperative behavior $p_{gi}=\frac{k_{g}}{K}p(gi|g)+\frac{K-k_{g}}{K}p(gi|g^{\prime}\neq g)$ is small if most interactions are external, as illustrated in Fig.3A, but cooperation will now be stabilized within each group. We note in Fig.3B that this need not be the final situation. Coordination and cooperation can continue to spread across groups selectively, as larger sets of group markers form new collective identities associated with in-group behavior through the mechanisms described in Table 1D. In this way, pedigree relations may be the most natural basis for in-group formation [22], but mutualistic strategic interactions can also spread if they become associated with recognizable characteristics and beneficial phenotypic behaviors.
@@ Discussion
We have demonstrated how to resolve several conceptual and practical problems in evolutionary theory by modeling fitness as an explicit process of adaptation to statistical environments. This result arises from interpreting population type frequencies as probabilities and assuming that adaptation is an optimal encoding of environmental evidence through Bayesian inference–a standard mapping in evolutionary dynamics [47, 46].
@@
The approach and results of this paper are part of a broad effort in the literature to make information more central in biology and evolution [65, 49, 19, 18, 45, 47, 14]. To this end, many authors have identified information theory as a natural language for biological phenomena ranging from the evolution of genome organization[65, 66], the emergence of structure in ecological networks[67], and fitness as it relates to environmental stochasticity [12, 14]. By formalizing the environment as a learnable statistical model, our approach replaces the circular notion of fitness as reproductive success by a well-defined statistical quantity that is optimized through recursive inference and characterized by information. In introducing this mechanism, we open up several directions for future research.
@@
First, defining fitness as a likelihood can guide both field observations and laboratory studies of evolutionary dynamics. This approach emphasizes identifying environmental states $e$ that yield fitness payoffs via measurement of $A$, and their statistical relations to specific phenotypes, $s$. This approach reduces (or compresses, in probability) the dimensionality of large spaces of environmental states to those with fitness consequences when conditioned on given phenotypes. Statistically independent factors can be identified as additive (in multiple variable linear regression) in the log-fitness (average growth rates). Practical applications in empirical settings also require identifying sources of variation not associated with selection. In standard evolutionary theory, mutations deviate population dynamics from pure Bayesian updating. Although fitness is still updated through the likelihood function, the dynamics now reflect a mutation-selection balance. Non-selective forces such as genetic drift and gene flow prevent the complete optimization of information transfer between the population and the environment, leading to incomplete adaptation and fixation of dominant types. These factors introduce parameters, such as the magnitude of fluctuations, to the analysis which will typically increase the evolvability of populations and enrich phase-space possibilities for evolutionary transitions.
@@
Many of these possibilities will lead to increased group heterogeneity. Our analysis of dyadic interactions shows that clusters of related phenotypes can promote cooperative behavior, introducing an information signal that can drive selection as argued in the general problem of major evolutionary transitions [68]. This approach generalizes beyond pairwise interactions, such as complex divisions of labor. This allows collective interactions to be analyzed via the structure of conditional (in)dependence of states, as established in probabilistic graphical models [69]. These results strengthen the intersection of evolutionary biology and statistical learning theory by integrating evolutionary dynamics with causal inference, information theory, and statistical learning, thereby paving the way for interdisciplinary applications in artificial intelligence.

### v2

@@ Results
The quadratic form of the log-fitness in $p$ is familiar and simple to analyze. A distinction emerges depending on the relative magnitude of $b$ and $c$, Fig. 2A. When $b>c$ (or $WP>L^{2}$) the situation is simple because the log-fitness is always a growing function of $p$, with a single maximum at $p=1$ corresponding to full cooperation. Thus, whatever the initial state of the population, cooperation will always emerge as the dominant strategy and will eventually fixate. Relative to the statistically independent case, cooperation now emerges from the fact that the population is effectively playing itself, so that increases in mutual exploitation lead to overall lower fitness. The situation is more interesting when $c>b$, because this leads to a fitness minimum at $p_{*}=\frac{c-b}{d}<1$. The existence of this fitness minimum creates a local fitness maximum at $p=0$ (mutual defection), which is locally stable. In these circumstances, sufficient cooperation $p>p_{*}$ must be present for the population to converge to the global fitness maximum of full cooperation ($p=1$), Figure 2A. This corresponds to a simple condition for cooperation to emerge, related to Hamilton’s rule [55]. From Eq. 15, we identify benefits of cooperation as $B=b+dp$ and costs with $C=c$. Then $B>C$ is a condition for full cooperation to emerge, which is equivalent to $p>p_{*}$. The population dynamics around $p=p_{*}$ is characterized by a first-order transition, a "tipping point," because there is a valley of fitness to be crossed to reach the global maximum. This can be done locally, over space or on networks, as many authors have noted based on numerical experiments [27, 56, 38]. We will return to this point below.
@@
Third, the situation is more interesting and more complex when the agents’ actions become statistically dependent. In this case, there are three parameters at play in the joint distribution: four joint probabilities minus the normalization. We can parameterize this space in analogy to task selection as $p_{2}=f_{1|0}(1-p_{1})+f_{1|1}p_{1}$ and $f(s_{1}=1,s_{2}=1)=f_{1|1}p_{1}$, with the shorthand $f_{1|1}=f(s_{2}=1|s_{1}=1)$, $f_{1|0}=f(s_{2}=1|s_{1}=0)$. Note that $r=f_{1|1}-f_{1|0}$ is the regression [10, 22] of $p_{2}$ on $p_{1}$, which vanishes when the two agents are statistically independent. Thus, $r$ is a measure of how (co-)related the two agents’ phenotypes are. The log-fitness becomes
@@
Because this is linear in $p_{1}$ the nature of the maximum depends only on the sign of the term in the square brackets. This is a second version of Hamilton’s rule closer to the way Queller [22] expressed it. If we identify the benefits of interaction $B=b$, and costs $C=c$ we now obtain $rB>C$ as the condition for cooperation to emerge. Adding the $d$ term facilitates the emergence of cooperation via what Queller called a synergism coefficient, as $rB+df_{1|1}>C$. This transition has a geometric interpretation in the reciprocity-exploitation ($f_{1|1},f_{1|0}$) plane. We can write the cooperation condition as a straight line, $y=f_{1|1}$, $x=f_{1|0}$, $y=y_{0}+\beta x$, with $y_{0}=\frac{c}{b+d}$, $\beta=\frac{b}{b+d}$ , as shown in Figure 2C. Cooperation, $p_{1}\rightarrow 1$, emerges for values above this line. To be above the line requires controlling exploitation, optimized when $f_{1|0}=f(s_{2}=1|s_{1}=0)\rightarrow 0$, which implies that punishment or retaliation is strong, $f(s_{2}=0|s_{1}=0)\rightarrow 1$. It also requires reciprocity meaning that $f_{1|1}=f(s_{2}=1|s_{1}=1)\rightarrow 1$, so that one becomes safe from exploitation, $f(s_{2}=0|s_{1}=1)\rightarrow 0$. Note that the first of these conditions may take the population towards mutual defection [57], which is suboptimal. The second is therefore necessary to reach the highest fitness globally. Below the line, cooperation cannot emerge. These outcomes agree with established results on the effects of reciprocity, punishment and rewards on cooperation [58, 59]. That is, both punishment and reciprocity promote the establishment of cooperation but alone punishment does not suffice [60, 57]. The information in the population is maximal as agents coordinate but the amount of information at stake converges to zero at full cooperation because then there is no uncertainty left to explain.
@@
Figure 2: Phase diagram for the Prisoner’s Dilemma and the environmentally mixed game. A: Log-fitness versus cooperation probability $p$ for the well-mixed population. Note the fitness valley for $b<c$. Conditional dependent case conditions for cooperation. B: Probability of reciprocity $f_{1|1}$ versus environment, $q$. C: Probability of reciprocity versus exploitation $f_{1|0}$, for various $q$. Cooperation emerges in each case for parameters above the lines. See also Fig. 1 of the supplement
@@
$\ln w_{s_{1}}=a-c(e)s_{1}+b(e)s_{2}+ds_{1}s_{2},$ (17)
with $c(e)=c(1-e)-be$ , $b(e)=b(1-e)-ce$. Costs and benefits swap between the PD and the ToC, for which $\ln w_{s_{1}}=a+bs_{1}-cs_{2}+ds_{1}s_{2}$. If the two agents act independently, then the agent should always cooperate in the ToC because both $b,d>0$. If the two agents share the same probability $p_{1}=p_{2}=p$, the log-fitness is independent of $e$ because there is a symmetry between the two agents so that the benefits and costs of cooperation are fully socialized. Thus, in the fully mixed population, changes in the environment have no effect on the average population dynamics and the phase diagram is the same as in Fig.2B.
When the two agents coordinate, the situation becomes more interesting. If agents can tell which game they are playing at each time, their probability of cooperation will simply follow from the selected strategy for each game, weighed by their relative probability, $q$. Alternatively, we consider when agents act conditionally on each other, independently of the state of the environment. Then we obtain the same condition for cooperation to emerge as in PD, but with costs and benefits that depend on the environment as in Eq. 17. Fig. 2B-C shows that when $q=0$, we obtain the PD condition (black line) in Fig. 2. As $q$ increases, both the ordinate at the origin and the slope decrease. At $q=1/2$ the slope is zero. Eventually, at $q=2/3$, ToC is sufficiently dominant and cooperation emerges unconditionally.
This analysis is general and can be obtained for any combination of games, leading to complex and interesting phase spaces, Fig. 2. As a result, we obtain transitions between different (optimal) population states as functions of environment probabilities, which play the role of control parameters analogous to temperature or pressure in physical systems. We have therefore shown how mixing the prisoner’s dilemma with a cooperative game increases the selection for cooperation in non-trivial ways. This is likely a more realistic situation as organisms, and especially people, have a diversity of modes of interaction, some competitive and some cooperative. Thus, the presence of explicitly cooperative channels some of the time helps unstable cooperation to emerge all of the time.
@@
The probabilistic interpretation of fitness extends naturally to structured populations. Within a population, group identities specify additional layers of organization that condition phenotypes and therefore may contribute to individual selection.
We write the fitness of individual $i$ in group $g$ as $w_{gi}(g^{\prime},j)=Ap(\{s_{g^{\prime}j}\}|s_{gi})$, where $\{s_{g^{\prime}j}\}$ denotes the joint state of all neighbors $j,g^{\prime}$. Typically, these neighbors are not statistically independent, leading to synergistic interactions [23, 61]. This is important for finding clusters of states (genes, phenotypes) that collectively encapsulate adaptive functions [62].
However, the most common situation arises when interactions are dyadic (between pairs only), so that the game occurs on a simple graph [63, 64, 35, 56]. In the log-fitness, this approximation is obtained when all the $s_{jg^{\prime}}$ are conditionally independent given $s_{ig}$ such that $w_{gi}=\Pi_{j,g^{\prime}}Ap(s_{g^{\prime}j}|s_{gi})$, resulting in
$\ln w_{g,i}=\ln A+\sum_{g^{\prime},j}\ln p(s_{g^{\prime}j}|s_{gi}).$ (18)
With these assumptions, we can write the log-fitness to be identical to Eq. 14, with the second player replaced by the average over all connections, $s_{2}=1/K\sum_{g^{\prime},j}{s_{g^{\prime}j}}$. All our previous results still apply but require attention to group dependencies, see Supplementary Information for details.
@@
In particular, the agent should still defect unconditionally when it is independent from the other agents. When their probabilities coincide, in the well-mixed population $p_{gi}=p_{2}=p$, Eq. 15 still applies and a phase transition occurs at $p=p_{*}$, if $b<c$. When the states are not independent, the log-fitness is written compactly in terms of contributions from each interaction, weighed by population size $1/K$,
@@
This equation is similar to what we obtained above, but with population definitions of $r$, $f_{1|1g},f_{1|0g}$, where the additional $g^{\prime},g$ labels denote conditional dependence across groups; the absence of the label means average across groups. The interesting possibility is that the agents’ behaviors are now different within and across groups in the sense that reciprocity $f_{1g^{\prime}|1g}<f_{1g|1g}$, and exploitation $f_{1g^{\prime}|0g}>f_{1g|0g}$, for $g^{\prime}\neq g$. Treating out group interactions as independent leads to
$\frac{1}{K}\ln w_{gi}=$ $a+\underbrace{\frac{k_{g}}{K}\left(br_{g}+df_{1g|1g}-c\right)}_{\rm in-group}p_{gi}$ (20)
$+\underbrace{\sum_{g^{\prime}\neq g}\frac{k_{g^{\prime}}}{K}\big{[}\left(dp_{g^{\prime}}-c\right)p_{gi}+bp_{g^{\prime}}\big{]}}_{\rm out-groups}.$ (21)
@@
This collapse of cooperation under external pressure can be averted in two ways. First, the agent may cease to interact with out-group members so that $k_{g^{\prime}\neq g}=0$, but this foregoes fitness benefits from the possible altruism of a large population. Fig.3A shows when cooperation emerges, solving $rB>C$ for $k_{g}/K$ as a function of $p_{\bar{g}}$. As before, the contours define a boundary, above which more cooperation increases fitness. We also plot this solution in Fig. 3B for $f_{1g|1g}$ and $p_{\bar{g}}$, which is comparable to Fig. 2B when $q=0$.
Second, cooperation or defection could conditionally depend on the group identity of the interaction partner. Then the in-group and out-group probabilities of cooperation $p_{gi|g^{\prime}}$ evolve in opposite directions so that $p(gi|g)\rightarrow 1$, and $p(gi|g^{\prime}\neq g)\rightarrow 0$. This may still mean that the average cooperative behavior $p_{gi}=\frac{k_{g}}{K}p(gi|g)+\frac{K-k_{g}}{K}p(gi|g^{\prime}\neq g)$ is small if most interactions are external, as illustrated in Fig.3A, but cooperation will now be stabilized within each group. We note in Fig.3B that this need not be the final situation. Coordination and cooperation can continue to spread across groups selectively, as larger sets of group markers form new collective identities associated with in-group behavior through the mechanisms described in Table 1D. In this way, pedigree relations (kin selection) may often be the most natural basis for in-group formation [22], but mutualistic interactions can also spread if they become associated with recognizable collective signals and beneficial phenotypic behaviors.
@@ Discussion
We demonstrated that defining fitness quantitatively as an explicit process of adaptation to statistical environments resolves several fundamental conceptual and practical problems in evolutionary theory. This re-definition arises necessarily from interpreting population type frequencies as probabilities and assuming that adaptation is an optimal encoding of environmental evidence through Bayesian learning–an argument advanced recently by several authors [47, 46].
@@
The approach and results of this paper are part of a broad effort in the literature to make information more central in biology and evolution [65, 49, 18, 17, 45, 47, 13]. To this end, many authors have identified information theory as a natural language for biological phenomena ranging from the evolution of genome organization [65, 66], the emergence of structure in ecological networks [67], and fitness as it relates to environmental stochasticity [11, 13]. By formalizing the environment in terms of learnable statistics associated with payoffs to growth, our approach replaces the circular notion of fitness as reproductive success by a well-defined statistical quantity that is optimized through recursive inference and characterized by information.
@@
In introducing this connection, we hope to open up several directions for future research. First, defining fitness as a likelihood can guide both field observations and laboratory studies of evolutionary dynamics. This approach emphasizes identifying environmental states $e$ that yield fitness payoffs via measurement of $A$, and their statistical relations to specific phenotypes, $s$. This approach reduces (or compresses, in probability) the dimensionality of large spaces of environmental states to those with fitness consequences when conditioned on given phenotypes. Statistically independent factors can be identified as additive (in multiple variable linear regression) in the log-fitness (average growth rates). Practical applications in empirical settings also require identifying sources of variation not associated with selection. As in standard evolutionary theory, mutations deviate population dynamics from pure Bayesian updating. Although fitness is still updated through the likelihood function, the dynamics now reflect a mutation-selection balance. Other non-selective forces such as genetic drift and gene flow, prevent the complete optimization of information transfer between the population and the environment, leading to incomplete adaptation and fixation of dominant types. These factors introduce parameters, such as the strength of inter-group interactions, to the analysis which may increase the evolvability of populations and enrich phase-space possibilities for evolutionary transitions [68].
@@
Many of these possibilities will lead to increased intra-group heterogeneity. Our analysis of dyadic interactions shows that clusters of related phenotypes can promote cooperative behavior, introducing an information signal that can drive selection as argued in the general problem of major evolutionary transitions [69]. This approach generalizes beyond pairwise interactions, such as complex divisions of labor. This allows collective interactions to be analyzed via the structure of conditional (in)dependence of states, as established in probabilistic graphical models [70]. These results strengthen the intersection of evolutionary biology and statistical learning theory by integrating evolutionary dynamics with causal inference, information theory, and recent progress in machine learning.